# ARFlow: Autogressive Flow with Hybrid Linear Attention

### v1

@@ Abstract
Flow models are effective at progressively generating realistic images, but they generally struggle to capture long-range dependencies during the generation process as they compress all the information from previous time steps into a single corrupted image. To address this limitation, we propose integrating autoregressive modeling—known for its excellence in modeling complex, high-dimensional joint probability distributions—into flow models. During training, at each step, we construct causally-ordered sequences by sampling multiple images from the same semantic category and applying different levels of noise, where images with higher noise levels serve as causal predecessors to those with lower noise levels. This design enables the model to learn broader category-level variations while maintaining proper causal relationships in the flow process. During generation, the model autoregressively conditions the previously generated images from earlier denoising steps, forming a contextual and coherent generation trajectory. Additionally, we design a customized hybrid linear attention mechanism tailored to our modeling approach to enhance computational efficiency. Our approach, termed ARFlow, under 400k training steps, achieves 14.08 FID scores on ImageNet at 128 × 128 without classifier-free guidance, reaching 4.34 FID with classifier-free guidance 1.5, significantly outperforming the previous flow-based model SiT’s 9.17 FID. Extensive ablation studies demonstrate the effectiveness of our modeling strategy and chunk-wise attention design.
@@ 1 Introduction
Furthermore, considering that ARFlow is expected to generate long sequences of image tokens—especially during autoregressive inference—we introduce a carefully designed linear attention mechanism as a replacement for self-attention, ensuring that the framework remains computationally feasible and scalable.
Specifically, since the autoregressive relationships in ARFlow are modeled at the image level rather than the patch level, we propose a tailored hybrid linear attention mechanism that groups tokens into chunks corresponding to different images.
By treating each image as a chunk, this mechanism implements full attention within chunks to model the generation of a single image, while applying causal masking only to inter-chunk attention to model image-wise relationships. This design preserves the necessary causality introduced by ARFlow and enables efficient parallel processing.
Extensive experiments on ImageNet at 128$\times$128 resolution are provided to demonstrate the effectiveness of ARFlow. For example, ARFlow achieves an FID of 4.34 with classifier-free guidance of 1.5, significantly outperforming the flow model SiT (FID 9.17) on 400k training steps. Moreover, our ablation studies further demonstrate three key strengths of our framework: 1) longer sequence lengths consistently enhance generation quality, with FID improving from 29.12 to 25.01 as sequences increase from 2 to 10 steps; 2) the cached state mechanism proves essential, as its removal results in a FID degradation from 25.46 to 65.33; and 3) our hybrid attention design can efficiently process long autoregressive sequences, rendering ARFlow’s generation speed comparable to SiT.
@@ 3.2 Flow models
$\mathbf{Z}_{t}=\alpha_{t}\mathbf{Z^{*}}+\sigma_{t}{\bm{\varepsilon}},$ (1)
@@
$\mathcal{L}_{\mathrm{v}}(\theta)$ $=\int_{0}^{T}\mathbb{E}[\|{\mathbf{v}}_{\theta}(\mathbf{Z}_{t},t)-\dot{\alpha}_{t}\mathbf{Z}^{*}-\dot{\sigma}_{t}{\bm{\varepsilon}}\|^{2}]{\mathrm{d}}t.$ (3)
@@ 3.3 Chunkwise linear attention
$\mathbf{S}_{[i+1]}$ $=\mathbf{S}_{[i]}+\sum_{t=iC+1}^{iC+C}\mathbf{k}_{t}\mathbf{v}_{t}^{\top}$ (4)
$=\mathbf{S}_{[i]}+\mathbf{K}_{[i+1]}^{\top}\mathbf{V}_{[i+1]}\in\mathbb{R}^{d\times d}.$ (5)
@@
$\mathbf{O}_{[i+1]}$ $=\underbrace{\mathbf{Q}_{[i+1]}\mathbf{S}_{[i]}}_{\text{inter-chunk}}$
$+\underbrace{\left((\mathbf{Q}_{[i+1]}\mathbf{K}_{[i+1]}^{\top})\odot\mathbf{M}\right)\mathbf{V}_{[i+1]}}_{\text{intra-chunk}}\in\mathbb{R}^{C\times d}$ (6)
@@ 4.1 ARFlow
Second, we transform these images into latent representations $\mathbf{Z}_{n}^{*}$ through our encoder (Equation 1), then perturb them with independently sampled noise $\bm{\varepsilon}_{n}$ according to the flow process (Equation 2). The independence of noise samples further enhances the diversity of the learning signal. Finally, we arrange the sequence based on the temporal ordering of $t_{n}$:
@@
$\mathcal{L}_{\theta}=\int_{0}^{T}\sum_{n=1}^{N}\mathbb{E}[\|v_{\theta}(\mathbf{Z}_{t_{n}}^{n},...,\mathbf{Z}_{t_{0}}^{0},t)-\dot{\alpha}_{t}\mathbf{Z}_{n}^{*}-\dot{\sigma}_{t}\bm{\varepsilon}_{n}\|^{2}]dt$ (8)
@@ 4.2 Hybrid linear attention
$g_{t}=\operatorname{sigmoid}\left(\mathbf{W}_{\gamma}\bm{x}_{t}\right)^{1/\tau}\in(0,1)$ (10)
where $\bm{x}_{t}\in\mathbf{X}$ is the input for each step, and $\tau=16$ is the temperature used for regulating the decaying factor to $1$ [39, 32] for retaining more historical information. Within each chunk $(i+1)$, we compute the geometric mean of the gating values to obtain the chunk-level decay:
@@ 5 Experiment Setup
We train a class-conditional latent model on the ImageNet dataset [6] $128\times 128$, a highly-competitive generative modeling benchmark. Following the original DiT [22], we initialize the final linear layer with zeros and use standard weight initialization techniques from ViT. All models are trained using the AdamW optimizer with a constant learning rate of $1\times 10^{-4}$, no weight decay, and a batch size of 256. Unless otherwise noted, the sequence length is set to 5. We maintain an exponential moving average (EMA) of model weights with a decay rate of 0.9999, and all reported results use the EMA model. The same training hyperparameters are used across all models, including the SiT [20] baselines.
@@
6 Experiments
@@ 6.1 Comparison with SiT
Table 1: Performance comparison between ARFlow and SiT models on ImageNet 128×128 with classifier-free guidance (CFG) = 1.0 and 1.5, across different model scales (S/2, B/2, L/2, XL/2). Metrics include Inception Score (IS), Fréchet Inception Distance (FID), saturated FID (sFID), Precision, and Recall.
[…]
SiT S/2 1.0 18.45 49.21 9.29 0.3764 0.5847
ARFlow S/2 1.0 22.15 41.56 8.47 0.4184 0.6060
SiT B/2 1.0 25.69 35.64 7.75 0.4539 0.6231
ARFlow B/2 1.0 36.73 25.46 7.13 0.5185 0.6489
SiT L/2 1.0 35.68 25.39 6.99 0.5211 0.6488
ARFlow L/2 1.0 55.03 15.91 6.43 0.5866 0.6651
SiT XL/2 1.0 37.59 24.02 6.92 0.5277 0.6481
ARFlow XL/2 1.0 60.78 14.08 6.46 0.5989 0.6705
SiT S/2 1.5 31.84 31.13 8.05 0.4902 0.5530
ARFlow S/2 1.5 41.48 23.95 7.56 0.5469 0.5779
SiT B/2 1.5 53.32 17.89 7.11 0.5984 0.5797
ARFlow B/2 1.5 84.18 10.16 6.77 0.6607 0.5908
SiT L/2 1.5 79.90 10.05 6.78 0.6780 0.5759
ARFlow L/2 1.5 130.12 4.92 6.42 0.7239 0.6020
SiT XL/2 1.5 85.10 9.17 6.83 0.6876 0.5755
ARFlow XL/2 1.5 140.07 4.34 6.38 0.7303 0.6106
@@
We compare ARFlow with SiT on ImageNet $128\times 128$ across different model scales and classifier-free guidance (CFG) settings in Tab. 1. ARFlow consistently outperforms SiT across all model scales and metrics. Without classifier guidance (CFG=1.0), ARFlow-XL/2 achieves FID 14.08 and IS 60.78, substantially improving upon SiT-XL/2’s 24.02 FID and 37.59 IS. With CFG=1.5(the improvements are more pronounced - ARFlow-XL/2 reaches 4.34 FID and 140.07 IS, marking a 52.67% relative FID improvement over SiT-XL/2. Enhanced Precision/Recall metrics demonstrate improvements in both sample quality and diversity.
Figure 3: Visual comparison of training loss between different scale SiT and ARFlow.
@@
As shows in Fig. 3, we visualize the training loss of SIT and ARFlow for different scales, and we can find that for the same model size, ARFlow’s loss is lower than that of the original SiT, and in particular, ARFlow-B5’s loss is close to SiT-XL’s loss.
Fig. 4 visualizes the sampling capabilities of ARFlow at 400K training steps, across different model sizes, using the same starting noise and seed. All images are generated with CFG=4.0. The comparison shows a clear improvement in image quality and fidelity as the model size increases.
Figure 4: Visual comparison of samples between different size ARFlow with CFG=4.0, demonstrating progressive improvement in sample quality as model size increase.
@@ 6.2 Ablation Study
Table 2: Impact of sequence length on ARFlow-B/2 model performance with CFG = 1.0.
@@
1 1.0 5.23 141.14 50.42 0.1021 0.0730
2 1.0 31.62 29.12 7.16 0.4959 0.6397
3 1.0 33.22 27.66 7.15 0.5077 0.6469
4 1.0 35.92 25.76 7.08 0.5173 0.6550
5 1.0 36.73 25.46 7.13 0.5185 0.6489
10 1.0 37.96 25.01 7.16 0.5189 0.6543
1 1.5 5.50 135.71 47.13 0.1067 0.0251
2 1.5 68.37 13.00 6.86 0.6429 0.5734
3 1.5 73.14 12.02 6.79 0.6472 0.5796
4 1.5 81.01 10.57 6.65 0.6581 0.5903
5 1.5 84.18 10.16 6.77 0.6607 0.5908
10 1.5 86.43 9.97 6.51 0.6629 0.5966
We first investigate the impact of sequence length on model performance in Tab. 2. With CFG=1.0, increasing the sequence length from 2 to 5 steps steadily improves the FID from 29.12 to 25.46. This trend continues up to 10 steps, reaching an FID of 25.01. Similar improvements are observed with CFG=1.5, where the FID decreases from 13.00 (2 steps) to 10.16 (5 steps) and further to 9.97 (10 steps). It is worth emphasizing that ARFlow training with sequence length of 1 cannot perform autoregressive flow generation, as it inherently reduces to the conventional flow model due to the absence of historical context required for autoregressive conditioning.
Figure 5: Visual comparison of training loss between Original SiT and ARFlow in different training sequence length, demonstrating progressive improvement as sequence length increases.
@@
Fig. 5 illustrates the progressive improvement in model performance as the sequence length of ARFlow increases. The training loss curves demonstrate a clear trend: longer sequence lengths consistently result in lower training loss. Starting with a sequence length of 1, which has a similar loss to the original SiT model, each increase in sequence length leads to a noticeable reduction in training loss, with the sequence length of 10 achieving the lowest loss.
We also visualize samples from the ARFlow Base model trained with different sequence lengths in Fig. 6. The results indicate that the model’s performance improves as the sequence length increases.
@@
CFG Scale
Table 4: Analysis of classifier-free guidance (CFG) scale on ARFlow-B/2 model performance.
[…]
1.0 36.73 25.46 7.13 0.5185 0.6489
1.5 84.18 10.16 6.77 0.6607 0.5908
2.0 143.36 5.33 7.42 0.7596 0.5154
2.5 195.91 5.47 8.76 0.8197 0.4490
3.0 237.44 7.26 10.30 0.8542 0.3917
3.5 269.50 9.43 11.84 0.8761 0.3354
4.0 292.18 11.44 13.13 0.8899 0.2855
4.5 309.58 13.18 13.33 0.8977 0.2548
5.0 321.76 14.65 15.36 0.9022 0.2247
The classifier-free guidance strength reveals interesting trade-offs in generation behavior. The FID improves substantially from CFG=1.0 (25.46) to CFG=2.0 (5.33). However, further increasing CFG leads to a quality-diversity trade-off: while precision continues to improve (from 0.7596 at CFG=2.0 to 0.9022 at CFG=5.0), recall decreases significantly (from 0.5154 to 0.2247). This suggests that stronger guidance produces higher quality samples but at the cost of reduced diversity.
@@
w 1.0 36.73 25.46 7.13 0.5185 0.6489
wo 1.0 13.69 65.33 17.03 0.2517 0.4650
w 1.5 84.18 10.16 6.77 0.6607 0.5908
wo 1.5 18.53 49.76 16.72 0.3671 0.4581
Tab. 5 presents an ablation study that examines the importance of cached states in our hybrid linear attention. We compare our full model with a variant where the hidden state transfer between chunks is removed, effectively eliminating the temporal memory mechanism. The results clearly demonstrate the crucial role of cached states in maintaining generation quality.

### v2

# ARFlow: Autoregressive Flow with Hybrid Linear Attention

@@ Abstract
Flow models are effective at progressively generating realistic images, but they generally struggle to capture long-range dependencies during the generation process as they compress all the information from previous time steps into a single corrupted image. To address this limitation, we propose integrating autoregressive modeling—known for its excellence in modeling complex, high-dimensional joint probability distributions—into flow models. During training, at each step, we construct causally-ordered sequences by sampling multiple images from the same semantic category and applying different levels of noise, where images with higher noise levels serve as causal predecessors to those with lower noise levels. This design enables the model to learn broader category-level variations while maintaining proper causal relationships in the flow process. During generation, the model autoregressively conditions the previously generated images from earlier denoising steps, forming a contextual and coherent generation trajectory. Additionally, we design a customized hybrid linear attention mechanism tailored to our modeling approach to enhance computational efficiency. Our approach, termed ARFlow, achieves 6.63 FID scores on ImageNet at 256 × 256 without classifier-free guidance, reaching 1.96 FID with classifier-free guidance 1.5, outperforming the previous flow-based model SiT’s 2.06 FID. Extensive ablation studies demonstrate the effectiveness of our modeling strategy and chunk-wise attention design.
@@ 1 Introduction
Furthermore, considering that ARFlow is expected to generate long sequences of image tokens—especially during autoregressive inference—we introduce a carefully designed linear attention mechanism as a replacement for self-attention, ensuring that the framework remains computationally feasible and scalable. Specifically, since the autoregressive relationships in ARFlow are modeled at the image level rather than the patch level, we propose a tailored hybrid linear attention mechanism that groups tokens into chunks corresponding to different images. By treating each image as a chunk, this mechanism implements full attention within chunks to model the generation of a single image, while applying causal masking only to inter-chunk attention to model image-wise relationships. This design preserves the necessary causality introduced by ARFlow and enables efficient parallel processing.
Extensive experiments on ImageNet are provided to demonstrate the effectiveness of ARFlow. For example, ARFlow achieves an FID of 4.34 with classifier-free guidance of 1.5, significantly outperforming the flow model SiT (FID 9.17) on 400k training steps on resolution of $128\times 128$, and achieves FID of 1.96 outperforming the SiT(2.06) on 7M training steps on resolution of $256\times 256$. Moreover, our ablation studies further demonstrate three key strengths of our framework: 1) longer sequence lengths consistently enhance generation quality, with FID improving from 29.12 to 25.01 as sequences increase from 2 to 10 steps; 2) the cached state mechanism proves essential, as its removal results in a FID degradation from 25.46 to 65.33; and 3) our hybrid attention design can efficiently process long autoregressive sequences, rendering ARFlow’s generation speed comparable to SiT.
@@ 3.2 Flow models
$\mathbf{Z}_{t}=\alpha_{t}\mathbf{Z^{*}}+\sigma_{t}{\boldsymbol{\varepsilon}},$ (1)
@@
$\mathcal{L}_{\mathrm{v}}(\theta)$ $=\int_{0}^{T}\mathbb{E}[\|{\mathbf{v}}_{\theta}(\mathbf{Z}_{t},t)-\dot{\alpha}_{t}\mathbf{Z}^{*}-\dot{\sigma}_{t}{\boldsymbol{\varepsilon}}\|^{2}]{\mathrm{d}}t.$ (3)
@@ 3.3 Chunkwise linear attention
$\mathbf{S}_{[i+1]}$ $=\mathbf{S}_{[i]}+\sum_{t=iC+1}^{iC+C}\mathbf{k}_{t}\mathbf{v}_{t}^{\top}=\mathbf{S}_{[i]}+\mathbf{K}_{[i+1]}^{\top}\mathbf{V}_{[i+1]}\in\mathbb{R}^{d\times d}.$ (4)
@@
$\mathbf{O}_{[i+1]}$ $=\underbrace{\mathbf{Q}_{[i+1]}\mathbf{S}_{[i]}}_{\text{inter-chunk}}+\underbrace{\left((\mathbf{Q}_{[i+1]}\mathbf{K}_{[i+1]}^{\top})\odot\mathbf{M}\right)\mathbf{V}_{[i+1]}}_{\text{intra-chunk}}\in\mathbb{R}^{C\times d}$ (5)
@@ 4.1 ARFlow
Second, we transform these images into latent representations $\mathbf{Z}_{n}^{*}$ through our encoder (Equation 1), then perturb them with independently sampled noise $\boldsymbol{\varepsilon}_{n}$ according to the flow process (Equation 2). The independence of noise samples further enhances the diversity of the learning signal. Finally, we arrange the sequence based on the temporal ordering of $t_{n}$:
@@
$\mathcal{L}_{\theta}=\int_{0}^{T}\sum_{n=1}^{N}\mathbb{E}[\|v_{\theta}(\mathbf{Z}_{t_{n}}^{n},...,\mathbf{Z}_{t_{0}}^{0},t)-\dot{\alpha}_{t}\mathbf{Z}_{n}^{*}-\dot{\sigma}_{t}\boldsymbol{\varepsilon}_{n}\|^{2}]dt$ (7)
@@ 4.2 Hybrid linear attention
$g_{t}=\operatorname{sigmoid}\left(\mathbf{W}_{\gamma}\boldsymbol{x}_{t}\right)^{1/\tau}\in(0,1)$ (9)
where $\boldsymbol{x}_{t}\in\mathbf{X}$ is the input for each step, and $\tau=16$ is the temperature used for regulating the decaying factor to $1$ [29, 34] for retaining more historical information. Within each chunk $(i+1)$, we compute the geometric mean of the gating values to obtain the chunk-level decay:
@@ 5.1 Experiment Setup
We train a class-conditional latent model on the ImageNet dataset [43] $128\times 128$ and $256\times 256$, a highly-competitive generative modeling benchmark. Following the original DiT [15], we initialize the final linear layer with zeros and use standard weight initialization techniques from ViT. All models are trained using the AdamW optimizer with a constant learning rate of $1\times 10^{-4}$, no weight decay, and a batch size of 256. Unless otherwise noted, the sequence length is set to 5. We maintain an exponential moving average (EMA) of model weights with a decay rate of 0.9999, and all reported results use the EMA model. The same training hyperparameters are used across all models, including the SiT [4] baselines.
@@
5.2 Comparison with DiT Backbone Models
Table 1 presents a performance comparison of ARFlow, SiT, and DiT models on the ImageNet $256\times 256$ dataset under varying classifier-free guidance (Cfg) conditions. Without classifier-free guidance, ARFlow consistently outperforms DiT and SiT, achieving notably superior results in IS (133.5), FID (6.63), and sFID (5.17). Under classifier-free guidance, ARFlow further extends its advantage, demonstrating the highest IS (288.4) alongside the lowest FID (1.96) and sFID (4.17). These results clearly indicate ARFlow’s superior capability in generating high-quality images compared to the baseline models, irrespective of guidance settings.
Table 1: Performance comparison between ARFlow, SiT and DiT on ImageNet 256×256 across model scales and classifier-free guidance values. Bold: The best.
[…]
DiT XL/2 7M w/o 121.5 9.62 6.85 0.67 0.67
SiT XL/2 7M w/o 131.7 8.61 6.32 0.68 0.67
ARFlow XL/2 7M w/o 133.5 6.63 5.17 0.68 0.65
DiT XL/2 7M w 278.2 2.27 4.60 0.83 0.57
SiT XL/2 7M w 270.3 2.06 4.49 0.83 0.59
ARFlow XL/2 7M w 288.4 1.96 4.17 0.81 0.61
@@ 5.3 Comparison with SiT
Table 2: Performance comparison between ARFlow and SiT on ImageNet 128×128 across model scales and classifier-free guidance values. Bold: The best.
(a) CFG = 1.0
Model Size IS FID Prec Rec
SiT S/2 18.45 49.21 0.3764 0.5847
ARFlow S/2 22.15 41.56 0.4184 0.6060
SiT B/2 25.69 35.64 0.4539 0.6231
ARFlow B/2 36.73 25.46 0.5185 0.6489
SiT L/2 35.68 25.39 0.5211 0.6488
ARFlow L/2 55.03 15.91 0.5866 0.6651
SiT XL/2 37.59 24.02 0.5277 0.6481
ARFlow XL/2 60.78 14.08 0.5989 0.6705
(b) CFG = 1.5
Model Size IS FID Prec Rec
SiT S/2 31.84 31.13 0.4902 0.5530
ARFlow S/2 41.48 23.95 0.5469 0.5779
SiT B/2 53.32 17.89 0.5984 0.5797
ARFlow B/2 84.18 10.16 0.6607 0.5908
SiT L/2 79.90 10.05 0.6780 0.5759
ARFlow L/2 130.12 4.92 0.7239 0.6020
SiT XL/2 85.10 9.17 0.6876 0.5755
ARFlow XL/2 140.07 4.34 0.7303 0.6106
@@
We compare ARFlow with SiT on ImageNet $128\times 128$ across different model scales and classifier-free guidance (CFG) settings in 2(b). ARFlow consistently outperforms SiT across all model scales and metrics. Without classifier guidance (CFG=1.0), ARFlow-XL/2 achieves FID 14.08 and IS 60.78, substantially improving upon SiT-XL/2’s 24.02 FID and 37.59 IS. With CFG=1.5(the improvements are more pronounced - ARFlow-XL/2 reaches 4.34 FID and 140.07 IS, marking a 52.67% relative FID improvement over SiT-XL/2. Enhanced Precision/Recall metrics demonstrate improvements in both sample quality and diversity.
@@
As shows in Figure 3(a), we visualize the training loss of SIT and ARFlow for different scales, and we can find that for the same model size, ARFlow’s loss is lower than that of the original SiT, and in particular, ARFlow-B5’s loss is close to SiT-XL’s loss.
(a) Training loss across different-scale models.
(b) Training loss under varying sequence lengths.
Figure 3: Comparison of training loss between SiT and ARFlow across scale and sequence length.
@@ 5.4 Ablation Study
We first investigate the impact of sequence length on model performance in Table 3(b). With CFG=1.0, increasing the sequence length from 2 to 5 steps steadily improves the FID from 29.12 to 25.46. This trend continues up to 10 steps, reaching an FID of 25.01. Similar improvements are observed with CFG=1.5, where the FID decreases from 13.00 (2 steps) to 10.16 (5 steps) and further to 9.97 (10 steps). It is worth emphasizing that ARFlow training with sequence length of 1 cannot perform autoregressive flow generation, as it inherently reduces to the conventional flow model due to the absence of historical context required for autoregressive conditioning.
Table 3: Impact of sequence length on ARFlow-B/2 performance under different classifier-free guidance (CFG) strengths.
@@
1 5.23 141.14 50.42 0.1021 0.0730
2 31.62 29.12 7.16 0.4959 0.6397
3 33.22 27.66 7.15 0.5077 0.6469
4 35.92 25.76 7.08 0.5173 0.6550
5 36.73 25.46 7.13 0.5185 0.6489
10 37.96 25.01 7.16 0.5189 0.6543
(a) CFG = 1.0
[…]
1 5.50 135.71 47.13 0.1067 0.0251
2 68.37 13.00 6.86 0.6429 0.5734
3 73.14 12.02 6.79 0.6472 0.5796
4 81.01 10.57 6.65 0.6581 0.5903
5 84.18 10.16 6.77 0.6607 0.5908
10 86.43 9.97 6.51 0.6629 0.5966
(b) CFG = 1.5
@@
Figure 3(b) illustrates the progressive improvement in model performance as the sequence length of ARFlow increases. The training loss curves demonstrate a clear trend: longer sequence lengths consistently result in lower training loss. Starting with a sequence length of 1, which has a similar loss to the original SiT model, each increase in sequence length leads to a noticeable reduction in training loss, with the sequence length of 10 achieving the lowest loss.
We also visualize samples from the ARFlow trained with different sequence lengths in Figure 4. The results indicate that the model’s performance improves as the sequence length increases.
CFG Scale
Table 4: Analysis of classifier-free guidance (CFG) scale on ARFlow-B/2 performance.
[…]
1.0 36.73 25.46 7.13 0.5185 0.6489
1.5 84.18 10.16 6.77 0.6607 0.5908
2.0 143.36 5.33 7.42 0.7596 0.5154
2.5 195.91 5.47 8.76 0.8197 0.4490
3.0 237.44 7.26 10.30 0.8542 0.3917
3.5 269.50 9.43 11.84 0.8761 0.3354
The classifier-free guidance strength reveals interesting trade-offs in generation behavior. The FID improves substantially from CFG=1.0 (25.46) to CFG=2.0 (5.33). However, further increasing CFG leads to a quality-diversity trade-off: while precision continues to improve (from 0.7596 at CFG=2.0 to 0.8761 at CFG=3.5), recall decreases significantly (from 0.5154 to 0.3354). This suggests that stronger guidance produces higher quality samples but at the cost of reduced diversity.
@@
w 1.0 36.73 25.46 0.5185 0.6489
wo 1.0 13.69 65.33 0.2517 0.4650
w 1.5 84.18 10.16 0.6607 0.5908
wo 1.5 18.53 49.76 0.3671 0.4581
Table 6 presents an ablation study that examines the importance of cached states in our hybrid linear attention. We compare our full model with a variant where the hidden state transfer between chunks is removed, effectively eliminating the temporal memory mechanism. The results clearly demonstrate the crucial role of cached states in maintaining generation quality.